8. The magnetic flux through a loop is given by Φm=0.6t^2-1.40t, where Φm is in webers and t is in seconds. Find the induced emf as a function of time. A) Φm=1.2t-1.4; B) Φm=-0.6t+1.4; C) Φm=-1.2t+1.4; D) Φm=1.2t-1.4;
Answer #1

Induced Emf is given by:

EMF = -N*d(\phi_m)/dt

\phi_m = 0.6*t^2 - 1.40*t

N = Number of loops = 1

d(\phi_m)/dt = d[0.6*t^2 - 1.40*t]/dt

d(\phi_m)/dt = 0.6*2*t - 1.40*1 = 1.2*t - 1.40

So,

EMF = -1*(1.2*t - 1.4)

EMF = -1.2*t + 1.4

Correct option is C.

Please Upvote.